OpenStudy (anonymous):

Subtract. Simplify, if possible (3-y)/(y-9)-(7y-9)/(9-y)

hero (hero):

(3-y)/(y-9) + (7y-9)/(y-9) = 6y-6/(y-9) = 6(y-1)/(y-9)